OpenStudy (anonymous):

Linear Algebra I: Let A be a 2x2 matrix and suppose that A has eigenvalues -1 and 3. Let I denote the 2x2 identity matrix. Prove that A² - 4A + 3 = 0

OpenStudy (anonymous):

My attempt: I tried working backwards to the eigenvalues (A-3)(A-1) = 0 but doesn't this imply eigen values are 3, 1 and not 3,-1? Any help would be appreciated

OpenStudy (jamesj):

(A-3)(A-1) = 0 isn't a valid matrix equation as for instance A - 3 doesn't make sense, a matrix minus a number. But even if you write (A-3I)(A-I) you're still not going to get exactly what you want. Yet you are right that there's a problem with this question. The sign of one of the eigenvalues is wrong or a sign in the quadratic is wrong.

OpenStudy (anonymous):

Okay so lets we assumed the eigenvalues were 1,3, how would be go about solving the problem then?

OpenStudy (jamesj):

A is similar to the diagonal matrix of eigenvalues. Suppose D = 3 0 0 1 Then there exists an invertible matrix P such that \[ A = P^{-1}DP \] Now given that, you can show A satisfies the quadratic.

OpenStudy (anonymous):

Ohhh so because of the eigenvalues we know the diagonal matrix D is. So if we assume A is dionilizable and there exist a invertible matrix P, then (P^-1 D P)² - 4(P^-1 D P) + 3I = 0, if we show this we are done, right?

OpenStudy (jamesj):

We don't have to assume A is diagonalizable. Given two distinct eigenvalues it's a theorem that A is diagonalizable. Note that \[ A^n = P^{-1}D^nP \] for all \( n \geq 1 \) and therefore given any polynomial p(A), \[ p(A) = P^{-1}p(D)P \]

OpenStudy (anonymous):

Write, i agree with that

OpenStudy (jamesj):

so if you show D satisfies your quadratic, it follows A satisfies it.

OpenStudy (anonymous):

ohhh I get it so we just gotta show D² - 4D + 3I = 0 is satisfied, then that proves A must be satisfied as well? right?

OpenStudy (jamesj):

Yes. Just write it out carefully.

OpenStudy (anonymous):

Got it, thank you very much :)